### v2

# Exact 3-D Channel Impulse Response Under Uniform Drift for Absorbing Spherical Receivers

Abstract
An exact channel impulse response (CIR) for the three-dimensional point-to-sphere absorbing channel under drift has remained unavailable due to symmetry breaking. This letter closes this gap by deriving an exact analytical CIR for a fully absorbing spherical receiver under uniform drift with arbitrary direction. By formulating the problem in terms of joint first-hitting time–location statistics and applying a Girsanov-based measure change, drift effects are isolated into an explicit multiplicative factor, yielding an exact series representation. The resulting CIR provides a rigorous reference model and enables efficient, noise-free evaluation of key channel metrics without relying on Monte Carlo simulations.
I Introduction
For three-dimensional molecular communication (MC) systems [3, 12, 5], the point-to-sphere absorbing channel has long served as a canonical reference model [7]. Its analytical tractability, however, relies critically on strong geometric symmetries that are immediately broken once drift is introduced.
In practical MC environments, drift is not a secondary effect but a fundamental transport mechanism, arising from background flow [15, 8], externally induced bias fields [11, 4], or spatial concentration gradients. Even when the drift field is spatially uniform, its presence breaks the radial symmetry that underlies classical no-drift formulations [17]. As a result, the first-hitting statistics depend not only on the arrival time but also on the angular location at which particles are absorbed.
Despite its physical relevance, an exact channel impulse response (CIR) for the point-to-sphere geometry under arbitrary drift direction has remained unavailable. Existing analytical results are restricted to zero-drift scenarios [17], aligned drift configurations [2], or reduced-dimensional approximations [15, 8], while more general settings are typically handled through Monte Carlo simulations or empirical fitting [7, 1]. This lack of an exact reference model has limited both physical interpretability and systematic channel analysis in drift-dominated regimes.
In this letter, we close this gap by deriving an exact analytical CIR for an absorbing spherical receiver [^1] In the molecular communication literature, a fully absorbing spherical receiver is widely used as a standard model for ligand–receptor binding or perfect capture at the receiver surface. in 3-D space under uniform drift with arbitrary direction. Rather than modifying existing no-drift expressions, we formulate the problem at the level of the joint first-hitting time and location statistics. By applying a measure-change argument based on the Girsanov theorem [14], the effect of drift is incorporated as an explicit reweighting of boundary hitting events, while preserving the analytical structure of the baseline no-drift solution.
The resulting CIR admits an exact series representation that remains valid for arbitrary drift orientations and strengths. This formulation provides a rigorous reference model for drifted point-to-sphere channels, resolving a long-standing symmetry-breaking regime that has previously been treated only approximately or through simulation [7, 13, 9, 6]. Beyond channel characterization, the availability of an exact CIR enables efficient and noise-free extraction of key system metrics, such as peak time and peak amplitude [17, 16], which are difficult to obtain reliably from particle-based Monte Carlo methods.
The main contributions of this letter are summarized as follows:
  • (i)
    An exact analytical series expression (1) for the 3-D point-to-sphere CIR under uniform drift with arbitrary direction.
  • (ii)
    A measure-change-based framework (see Fig. 2), via the Girsanov theorem, that isolates drift effects into an explicit multiplicative factor and admits extension to other molecular channel models.
TABLE I: Comparison of related analytical CIR results in MC literature. Note that in 1-D space, spherical and planar geometries coincide as the receiver reduces to a point boundary.
Dim. No Drift Uniform Drift
$d=1$ Lévy distribution [8, 15] Inverse Gaussian distribution [8, 15]
$d=3$ (Planar) Exact analytical CIR [10] Exact analytical CIR (arbitrary direction) [10]
$d=3$ (Spherical) Exact analytical CIR [17] No exact analytical CIR in prior work; [2] restricted to aligned drift
[…]
Figure 1: System model of the 3-D molecular communication channel. A point transmitter is located at $\mathbf{x}_{0}$, and a fully absorbing spherical receiver of radius $r$ is at the origin. The medium is subject to a uniform drift $\mathbf{v}$ forming an angle $\psi=\angle(\mathbf{v},\mathbf{x}_{0})$.
II System Model
We consider an unbounded 3-D fluid environment $\mathbb{R}^{3}$. The transmitter (Tx) is modeled as a point source located at $\mathbf{x}_{0}\in\mathbb{R}^{3}$, and the receiver (Rx) is a fully absorbing sphere of radius $r$ centered at the origin, denoted by $\mathcal{S}_{r}$, with $|\mathbf{x}_{0}|>r$. The system geometry is illustrated in Fig. 1.
Information particles (molecules) are released at time $t=0$. From a microscopic perspective, the trajectory of a particle $\mathbf{X}_{t}$ follows the Itô stochastic differential equation [14]
$d\mathbf{X}_{t}=\mathbf{u}(t)\,dt+\sigma\,d\mathbf{B}_{t},\quad\mathbf{X}_{0}=\mathbf{x}_{0},$ (1)
where $\mathbf{B}_{t}$ is a standard 3-D Brownian motion, $D$ is the diffusion coefficient, and $\sigma=\sqrt{2D}$.
While the proposed framework applies to general drift profiles, this letter focuses on the practically relevant case of constant uniform drift, $\mathbf{u}(t)\equiv\mathbf{v}\in\mathbb{R}^{3}$, forming an angle $\psi=\angle(\mathbf{v},\mathbf{x}_{0})$ with the Tx–Rx axis. The relative strength of advection and diffusion is characterized by the Péclet number $\mathrm{Pe}=|\mathbf{v}|r/D$. The first hitting time is defined as $T\triangleq\inf\{t>0:\ |\mathbf{X}_{t}|=r\},$ and the CIR is the probability density function of $T$, denoted by $f_{\mathrm{3D}}^{(\mathbf{v})}(t)$.
From a macroscopic viewpoint, the corresponding particle concentration satisfies the advection–diffusion equation with absorbing boundary conditions. However, for arbitrary drift directions, direct PDE-based approaches become analytically intractable, which motivates the stochastic measure-change formulation adopted in the next section.
Yilmaz et al. [17] 3-D No-Drift CIR (Marginal Time) Yin [18] Baseline 3-D Joint Time-Position The Drift-Factor Method Apply Girsanov Theorem (i.e., Radon-Nikodym Derivative) Proposed Theorem 1 Exact 3-D CIR with Arbitrary Drift Direction Insufficient (Lacks spatial info) Basis for Measure ChangeIntegration over Sphere
Figure 2: Derivation flow based on drift-factor method. We start from Yin’s joint time-position distribution [18] and apply the Girsanov theorem to derive the 3D CIR in analytical form.
III Derivation of the Drifted 3-D CIR
This section derives the exact CIR for a spherical absorbing receiver in 3-D space under a uniform drift with arbitrary direction. The overall derivation flow is summarized in Fig. 2. The key idea is to bridge the well-established no-drift case and the general drifted case through a measure-change argument.
III-A Motivation: Why the Joint Distribution Is Necessary
Existing analytical results for spherical receivers without drift typically rely on the marginal first-hitting-time distribution [17]. However, under drift, advection introduces a strong directional bias: the contribution of each hitting event depends not only on when the particle hits the receiver, but also on where it hits the spherical boundary. Consequently, the marginal hitting-time distribution alone is insufficient to characterize the drifted CIR.
To retain this spatial information, we start from the joint probability density of the first hitting time $T$ and the hitting location $\mathbf{X}_{T}$ on the spherical boundary $\mathcal{S}_{r}$, denoted by $f_{\mathrm{3D}}^{(\mathbf{0})}(t,\mathbf{y})$ for $\mathbf{y}\in\mathcal{S}_{r}$. An explicit analytical expression for this joint density in the no-drift case is available from probability literature [18] and serves as the baseline of our derivation.
III-B Drift-Factor Representation via Girsanov Theorem
To incorporate drift, we employ the Girsanov theorem [14] to perform a change of measure from a reference no-drift process (measure $\mathbb{Q}$) to the drifted process (measure $\mathbb{P}$). Evaluated at the hitting time, this change of measure introduces a multiplicative drift factor, which analytically relates the drifted joint density to the no-drift one through
$f_{\mathrm{3D}}^{(\mathbf{v})}(t,\mathbf{y})=\exp\!\left(\frac{\mathbf{v}\cdot(\mathbf{y}-\mathbf{x}_{0})}{\sigma^{2}}-\frac{|\mathbf{v}|^{2}t}{2\sigma^{2}}\right)f_{\mathrm{3D}}^{(\mathbf{0})}(t,\mathbf{y}).$ (2)
Physically, the drift factor reweights each hitting event according to the work done by the drift along the particle trajectory. Hitting locations aligned with the drift direction are exponentially favored, while those against the drift are suppressed. This representation isolates the effect of drift into a single, explicit weighting term (the drift factor) while preserving the analytical structure of the no-drift joint distribution.
III-C Baseline Joint Density for the No-Drift Case
For a Tx located outside the spherical Rx (i.e., $|\mathbf{x}_{0}|>r$), the joint density of the hitting time-location for a 3-D Brownian motion without drift admits the following series representation, adapted from [18, Theorem 3.2]:
$f_{\text{3D}}^{(\mathbf{0})}(t,\mathbf{y})$ $=\frac{1}{2\pi}\frac{r}{|\mathbf{x}_{0}|}\sum_{m=0}^{\infty}\Big(m+\tfrac{1}{2}\Big)P_{m}\!\big(\cos\angle(\mathbf{x}_{0},\mathbf{y})\big)$
$\times$ $\int_{0}^{\infty}\frac{\lambda\,Z_{m+\frac{1}{2}}(|\mathbf{x}_{0}|/r,\lambda r/\sigma)}{J_{m+\frac{1}{2}}^{2}(\lambda r/\sigma)+Y_{m+\frac{1}{2}}^{2}(\lambda r/\sigma)}\exp\!\left(-\tfrac{1}{2}\lambda^{2}t\right)\,d\lambda.$ (3)
where $P_{m}(\cdot)$ denotes the Legendre polynomial and $Z_{\nu}(a,b)=J_{\nu}(ab)Y_{\nu}(b)-J_{\nu}(b)Y_{\nu}(ab)$ denotes the standard cross-product of Bessel functions under normalized radial coordinates.
III-D CIR as Surface Marginalization
The drifted CIR is defined as the marginal distribution of the random hitting time $T$, obtained by integrating the drifted joint density over the spherical boundary:
$f_{\mathrm{3D}}^{(\mathbf{v})}(t)=\int_{\mathcal{S}_{r}}f_{\mathrm{3D}}^{(\mathbf{v})}(t,\mathbf{y})\,ds(\mathbf{y}).$ (4)
Substituting (2) into (4), the $\mathbf{y}$-independent term $\exp\!\left(-|\mathbf{v}|^{2}t/2\sigma^{2}-\mathbf{v}\cdot\mathbf{x}_{0}/\sigma^{2}\right)$ can be factored out. The remaining task is to evaluate surface integrals of the form $e^{\mathbf{v}\cdot\mathbf{y}/\sigma^{2}}$ multiplied by the angular modes in (III-C).
III-E Mode-Wise Surface Integration
To evaluate these integrals, we adapt the following identity from [18, Lemma 2.5] specialized to the 3-D setting.
Lemma 1 (3-D Surface Integral Identity).
For any vector $\mathbf{c}\in\mathbb{R}^{3}$,
$\int_{\mathcal{S}_{r}}e^{\mathbf{c}\cdot\mathbf{y}}P_{m}\!\big(\cos\angle(\mathbf{x}_{0},\mathbf{y})\big)\,ds(\mathbf{y})$
$\qquad=2(r\pi)^{\frac{3}{2}}\left(\frac{|\mathbf{c}|}{2}\right)^{-\frac{1}{2}}I_{m+\frac{1}{2}}(|\mathbf{c}|r)\,P_{m}\!\big(\cos\angle(\mathbf{c},\mathbf{x}_{0})\big),$ (5)
where $I_{\nu}(\cdot)$ denotes the modified Bessel function of the first kind.
Applying Lemma 1 with $\mathbf{c}=\mathbf{v}/\sigma^{2}$, the angular argument reduces to $\angle(\mathbf{v},\mathbf{x}_{0})=\psi$. Substituting (III-C) into (4), the CIR decomposes into a summation over angular eigenmodes $m$, where each term involves a single surface integral that can be evaluated explicitly using Lemma 1. Collecting all prefactors yields the exact 3-D CIR stated below in Theorem 1.
Theorem 1 (Exact 3-D CIR under uniform drift).
In $\mathbb{R}^{3}$, the CIR for a spherical absorbing receiver of radius $r$ under a uniform drift $\mathbf{v}$ is
$f_{\mathrm{3D}}^{(\mathbf{v})}(t)=$ $-\exp\!\left(-\frac{\mathbf{v}\!\cdot\!\mathbf{x}_{0}}{\sigma^{2}}-\frac{|\mathbf{v}|^{2}t}{2\sigma^{2}}\right)\frac{\sqrt{2}\,\sigma}{\sqrt{\pi}\,|\mathbf{v}|^{1/2}|\mathbf{x}_{0}|^{1/2}}$
$\times\sum_{m=0}^{\infty}\Big(m+\tfrac{1}{2}\Big)I_{m+\frac{1}{2}}\!\left(\frac{|\mathbf{v}|r}{\sigma^{2}}\right)P_{m}(\cos\psi)$
$\times\int_{0}^{\infty}\frac{\lambda\,Z_{m+\frac{1}{2}}(|\mathbf{x}_{0}|/r,\lambda r/\sigma)}{J_{m+\frac{1}{2}}^{2}(\lambda r/\sigma)+Y_{m+\frac{1}{2}}^{2}(\lambda r/\sigma)}e^{-\frac{1}{2}\lambda^{2}t}\,d\lambda,$ (6)
where $\psi=\angle(\mathbf{v},\mathbf{x}_{0})$, $\sigma^{2}=2D$, $I_{\nu}(\cdot)$ is the modified Bessel function of the first kind, and $P_{m}(\cdot)$ is the Legendre polynomial.
This exact representation (1) serves as the basis for the numerical evaluation and performance characterization presented in the next section.
(a) Drift towards Rx, lower speed
(b) Transverse drift, lower speed
(c) Drift away from Rx, lower speed
(d) Drift towards Rx, higher speed
(e) Transverse drift, higher speed
(f) Drift away from Rx, higher speed
Figure 3: Validation of analytical CIR against Monte Carlo simulations. The analytical curves correspond to $N_{\text{tx}}f_{\text{3D}}^{(\mathbf{v})}(t)\,\Delta t$, i.e., the expected number of absorbed molecules within a fixed time bin $\Delta t=5\times 10^{-5}\,\mathrm{s}$. Top Row: Moderate drift speed ($|\mathbf{v}|=5\,\mu\text{m/s}$). Bottom Row: Higher drift speed ($|\mathbf{v}|=10\,\mu\text{m/s}$). Columns correspond to positive ($\psi=180^{\circ}$), transverse ($\psi=90^{\circ}$), and negative ($\psi=0^{\circ}$) drift directions.
IV Numerical Results and Discussion
In this section, we validate the derived exact analytical CIR against particle-based Monte Carlo simulations and give a brief discussion on how the proposed formulation enables efficient characterization of two key channel metrics: pulse amplitude and pulse peak time. All numerical evaluations and simulations are performed using MATLAB (MathWorks cloud-based environment). To ensure consistency with established MCvD channel modeling literature [17], the default system parameters are chosen typically as follows: diffusion coefficient $D=80\,\mu\mathrm{m}^{2}/\mathrm{s}$, receiver radius $r=10\,\mu\mathrm{m}$, and initial distance $|\mathbf{x}_{0}|=20\,\mu\mathrm{m}$. We consider two representative drift scenarios: a moderate drift speed ($|\mathbf{v}|=5\,\mu\mathrm{m/s}$) and a higher drift speed ($|\mathbf{v}|=10\,\mu\mathrm{m/s}$).
For Monte Carlo validation shown in Fig. 3, the analytical CIR is discretized over a fixed time bin $\Delta t$ to enable a direct comparison with particle-based simulations. In Figs. 4 and 5, the peak time $t_{\mathrm{peak}}$ is obtained by maximizing the continuous-time analytical CIR, while the corresponding peak amplitude is reported in terms of the expected number of absorbed molecules per time bin, $N_{\mathrm{tx}}f^{(\mathbf{v})}_{\mathrm{3D}}(t_{\mathrm{peak}})\Delta t$, using the same $\Delta t$ as in Fig. 3 for consistent scaling.
IV-A Validation against Monte Carlo Simulations
Fig. 3 compares the analytical CIR (Eq. (1), truncated at $m=30$) with particle-based Monte Carlo simulations using $N_{\mathrm{tx}}=10^{6}$ molecules, for drift magnitudes $|\mathbf{v}|\in\{5,10\}\,\mu\mathrm{m/s}$ and drift angles $\psi\in\{0^{\circ},90^{\circ},180^{\circ}\}$. Excellent agreement is observed across all considered scenarios.
As shown in Fig. 3, increasing drift aligned with the Tx–Rx axis sharpens the received pulse, whereas drift directed away from the receiver suppresses the peak and elongates the tail. These trends are accurately captured by the analytical CIR. Any residual discrepancies are attributable to finite-particle fluctuations inherent in Monte Carlo simulations.
IV-B Convergence of the Analytical Series
Numerical evaluation of Theorem 1 requires truncation of the infinite series over the angular mode index $m$. From the structure of (1), the truncation order $M$ controls the angular resolution of the CIR through the Legendre modes $P_{m}(\cos\psi)$, while temporal smoothing is governed by the Gaussian factor $e^{-\lambda^{2}t/2}$ in the radial integral.
For moderate drift strengths, higher-order angular modes are increasingly suppressed by the modified Bessel term $I_{m+\frac{1}{2}}(|\mathbf{v}|r/\sigma^{2})$, and truncation at $M=30$ yields a relative error below $10^{-5}$ over most of the time domain for the parameters considered in this work. This accuracy is sufficient for Monte Carlo validation and for the extraction of peak metrics.
Convergence becomes slower only in the immediate vicinity of the first arrival (very small $t$), where temporal smoothing is weak and sharper angular features must be resolved, or under stronger drift conditions that induce pronounced angular anisotropy. In such cases, increasing the truncation order to $M\approx 50$ provides accurate resolution of the initial transient. For practical operating regimes, $M=30$ offers an effective accuracy–complexity trade-off.
IV-C Characterization of Peak Metrics
Monte Carlo estimation of peak metrics is computationally expensive because extensive averaging is needed to suppress stochastic fluctuations. In contrast, the analytical CIR provides a smooth objective function, enabling efficient numerical maximization. Compared with particle-based Monte Carlo estimation, this approach reduces the computational cost by several orders of magnitude, as it avoids repeated stochastic sampling and averaging.
Fig. 4 reports peak metrics versus drift magnitude. The peak absorption count is reported as the expected number of absorbed molecules within a fixed time bin $\Delta t$, whereas $t_{\mathrm{peak}}$ is obtained directly from the continuous-time analytical CIR. As $|\mathbf{v}|$ increases, the peak count grows under positive drift and is strongly suppressed under negative drift, while $t_{\mathrm{peak}}$ decreases monotonically.
The trends observed in Fig. 5 arise from the underlying geometry of the absorbing receiver rather than from normalization or parameter choices. Increasing the receiver radius enlarges the absorbing surface, thereby enhancing the instantaneous capture probability and leading to a higher peak absorption count. At the same time, a larger radius reduces the effective propagation distance $|\mathbf{x}_{0}|-r$, which naturally advances the peak time.
Figure 4: Impact of drift magnitude $|\mathbf{v}|$ on peak metrics for a fixed receiver radius $r=10~\mu\mathrm{m}$. Top: peak absorption count per time bin, computed as $N_{\mathrm{tx}}f^{(\mathbf{v})}_{\mathrm{3D}}(t_{\mathrm{peak}})\Delta t$ with $\Delta t=5\times 10^{-5}\,\mathrm{s}$ (same bin width as in Fig. 3 for consistent scaling). Bottom: corresponding peak time $t_{\mathrm{peak}}$, obtained by maximizing the continuous-time analytical CIR and thus independent of $\Delta t$. Positive ($\psi=180^{\circ}$), transverse ($\psi=90^{\circ}$), and negative ($\psi=0^{\circ}$) drift directions are shown.
Figure 5: Impact of receiver radius $r$ on peak metrics for a fixed drift magnitude $|\mathbf{v}|=10~\mu\mathrm{m/s}$. Top: peak absorption count per time bin, computed as $N_{\mathrm{tx}}f^{(\mathbf{v})}_{\mathrm{3D}}(t_{\mathrm{peak}})\Delta t$ with $\Delta t=5\times 10^{-5}\,\mathrm{s}$ (same bin width as in Fig. 3 for consistent scaling). Bottom: corresponding peak time $t_{\mathrm{peak}}$, obtained by maximizing the continuous-time analytical CIR and thus independent of $\Delta t$. Positive ($\psi=180^{\circ}$), transverse ($\psi=90^{\circ}$), and negative ($\psi=0^{\circ}$) drift directions are shown.
V Conclusion
This letter establishes an exact analytical characterization of the 3-D CIR for MCvD systems with a fully absorbing spherical receiver under uniform drift of arbitrary direction. By formulating the problem at the level of joint first-hitting time and location statistics and applying a measure-change argument, we derive an exact series representation of the CIR within the drift–diffusion model, without relying on geometric approximations [2] or dimensional reductions [8]. The analytical CIR is validated against particle-based Monte Carlo simulations over representative drift magnitudes and orientations, showing excellent agreement. Beyond validation, the exact formulation enables efficient and noise-free extraction of key channel metrics via direct numerical maximization with respect to the CIR curve with substantially lower computational cost than simulation-based approaches. Numerical results further clarify how drift orientation and receiver geometry jointly shape the received signal, capturing enhancement, suppression, and geometric scaling effects in a unified manner. From a system-level perspective, the proposed CIR provides a precise and computationally tractable channel model that can be potentially integrated into molecular MIMO detection and receiver design (e.g., [6, 1]), completing a previously missing exact reference model for drifted point-to-sphere molecular channels [7].
Appendix A Methodological Consistency Check
This appendix provides a brief consistency check for the proposed measure-change framework. It is not used as a reduction or a derivation step for the 3-D spherical model in the main text.
In one dimension, an absorbing boundary implies a deterministic point hitting location on the hitting event. Let the initial distance be $\ell>0$ and $\sigma^{2}=2D$. The no-drift first-hitting-time density is $f_{\mathrm{1D}}^{(0)}(t)=\frac{\ell}{\sqrt{4\pi Dt^{3}}}\exp\!\left(-\frac{\ell^{2}}{4Dt}\right),~t>0$, see [15]. Under constant drift $v$ (towards the boundary), the Girsanov factor evaluated at the hitting time reduces to $\exp\!\left(\frac{v\ell}{2D}-\frac{v^{2}t}{4D}\right)$, corresponding to (2) with $\mathbf{v}\!\cdot\!(\mathbf{y}-\mathbf{x}_{0})=v\ell$ in one dimension. Multiplying this factor with the no-drift density $f_{\mathrm{1D}}^{(0)}(t)$ yields
$f_{\mathrm{1D}}^{(v)}(t)$ $=\frac{\ell}{\sqrt{4\pi Dt^{3}}}\exp\!\left(-\frac{\ell^{2}}{4Dt}+\frac{v\ell}{2D}-\frac{v^{2}t}{4D}\right)$
$=\frac{\ell}{\sqrt{4\pi Dt^{3}}}\exp\!\left(-\frac{(\ell-vt)^{2}}{4Dt}\right),$ (7)
where the exponent is obtained by completing the square. This recovers the standard inverse Gaussian first-passage-time law (cf. [8, 15]), which corresponds to the 1-D uniform-drift case as presented in Table I.